### v1

@@ 4.2 Results
The results of our estimation are reported in Table 5. The first 3 columns show estimates of the structural parameters for links of the same type, while the remaining 3 columns are the estimates for links of different types.
Following the $32$ identified library types in Figure 1, we estimate a model with $K=32$ unobserved types and initialize the type allocation for our algorithm using InfoMap (Rosvall et al., 2009). [^22] As suggested in the literature on variational approximation (Wainwright and Jordan, 2008), we started the algorithm from different initial guesses of the type allocations but only report the results that achieved the highest lower bound.
We estimate a positive externality ($\gamma$), suggesting that a simple stochastic blockmodel is unable to account for the dependence of the links in our data. This result aligns with our strategic model, suggesting that links are highly interdependent. Essentially, when coders form connections, they (partially) consider the indirect impacts of their actions. However, this behavior leads to a network architecture that might not be as efficient as one where a central planner aims to maximize the overall utility of each participant. Indeed, the externality encourages coders to establish more connections than would be ideal from an aggregate utility-maximization perspective. This also has implications for contagion, as we explain in the next section.
The parameter $\alpha$ governs the density of the network and is negative both between and within types. This implies that—other things being equal— in equilibrium sub-networks of libraries of the same type tend to be more sparse than between-types. One interpretation is that $\alpha$ measures costs of forming links. By contrast, the parameter $\beta_{1}$ is positive both within and across types, indicating that coders are more likely to link to larger packages of the same and of different types.
Table 5: Parameter estimates and standard errors. Estimates obtained by setting $K=32$ types and running the algorithm for 300 iterations, after initializing with InfoMap.
@@
Est. Std. Err. z value Est. Std. Err. z value
Externality ($\gamma$) .$264$002.002.002.002 $121.996$
Edges ($\alpha$) $-12.070$065.065.065.065 $-186.323$ $-6.602$002.002.002.002 $-3843.154$
Size($\beta_{1}$) .$166$049.049.049.049 $3.411$$432$003.003.003.003 $167.775$
Popularity($\beta_{2}$) .$546$048.048.048.048 $11.326$$-.060$003.003.003.003 $-19.365$
Num.Forks($\beta_{3}$) $-1.012$050.050.050.050 $-2.332$ $-2.266$004.004.004.004 $-507.326$
Num.Contributors($\beta_{4}$) .$063$047.047.047.047 $1.327$ $-1.12$004.004.004.004 $-314.475$
Maturity($\beta_{5}$) $2.659$061.061.061.061 $43.751$$-.448$003.003.003.003 $-141.542$
Interestingly, Popularity  has a positive coefficient $\beta_{2}$ within types and a—much smaller—negative coefficient between types. The same is true for Num.Contributors, which is an alternative measure of popularity. Coders are, therefore, likely to link to more popular packages of the same type but less popular packages of other types. One interpretation consistent with these results is that coders use popular features of same-type packages to include in their own code but use infrastructure-style code from different-type packages, which is not necessarily very popular. This interpretation is also consistent with the coefficients $\beta_{4}$ and $\beta_{5}$ for Num.Contributors and Maturity, which are positive within types and negative between-type packages. Lastly, the coefficient $\beta_{3}$ for Num.Forks is negative both within and across projects, but not only significantly for between-type links. This could be a mechanical effect, though: If a project has more forks, there are more similar versions of the same package, meaning that a coder has more possibilities to link to it.
@@ 5 Contagious Vulnerabilities
We adapt these ideas to the contagion of vulnerabilities in software dependency graphs. The prevalence of vulnerabilities in software have led to the creation of the Common Vulnerabilities and Exposure (CVE) program and the National Vulnerability Database (NVD) by the National Institute of Standards and Technology, which is an agency of the United States Department of Commerce (NIST, 2024). A software package depending on a vulnerable package is potentially vulnerable itself. To see how, consider, for example, the Equifax data breach (Federal Trade Commission, 2024) in 2017, caused by a vulnerability in the popular open source web application framework Apache Struts 2.232323Apache Struts 2 did not properly validate the Content-Type header of incoming http requests when processing file uploads, which allowed attackers to create http requests that include malicious code. Due to the lack of content validation, this allowed attackers to execute arbitrary code on the web server, ultimately granting them full control of the server. For details see the official NIST announcement on CVE-2017-5638 (https://nvd.nist.gov/vuln/detail/CVE-2017-5638). Equifax did not itself develop or maintain Apache Struts 2, but instead used it as part of its own code base. The vulnerability allowed attackers to execute malicious code providing access to the server hosting the web application developed by Equifax using Apache Struts 2 and costing the company more than \$1.5 Billion to date. [^24] See https://www.bankinfosecurity.com/equifaxs-data-breach-costs-hit-14-billion-a-12473. Article accessed 27 January 2024.
@@ 6 Conclusion
The network of dependencies among software packages is an interesting laboratory to study network formation and externalities. Indeed, the creation of modern software gains efficiency by re-using existing libraries; on the other hand, dependencies expose new software packages to bugs and vulnerabilities from other libraries. This feature of the complex network of dependencies motivates the interest in understanding the incentives and equilibrium mechanisms driving the formation of such networks.
@@
We find evidence that coders create positive externalities for other coders when creating a link. This raises more questions about the formation of dependency graphs. We study how vulnerable the observed network is to vulnerability contagion and show that ensuring that even a relatively small number of packages is vulnerability free effectively curbs the extent of contagion. As a next step, we could study how changing the parameters of the estimated network formation model affects vulnerability contagion. The presence of an externality implies that maintainers may be inefficiently creating dependencies, thus increasing the density of the network and the probability of contagion. The extent of this risk and the correlated damage to the system is of paramount importance and we plan to explore it in future versions. Second, while we have focused on a stationary realization of the network, there are important dynamic considerations in the creation of these dependencies. While the modeling of forward-looking maintainers may be useful to develop intuition about intertemporal strategic incentives and motives, we leave this development to future work. Finally, we have focused on a single language, but the analysis can be extended to other languages as well, such as Java, C++ and others.
@@ Appendix A Computational details for estimation
where each firm’s type is i.i.d. multinomial, so the distribution $p_{\bm{\eta}}(\bm{z})$ is
@@ A.1.1 Variational EM algorithm with MM updates
Then under some conditions – namely, that the network is large and each block/type is not too large with respect to the network (Babkin et al., 2020; Schweinberger, 2020; Schweinberger and Stewart, 2020) – we have
$L\left(\bm{g},\bm{x},\bm{z};\bm{\alpha},\bm{\beta},\gamma,\bm{\eta}\right)\approx L_{0}\left(\bm{g},\bm{x},\bm{z};\bm{\alpha},\bm{\beta},\bm{\eta}\right)$ (11)
Variational methods for the stochastic blockmodel are relatively standard and involve estimating an approximating distribution $q_{\xi}(\bm{z})$ that minimizes the Kulback-Leibler divergence from the true likelihood. This can be achieved in several ways, but usually, the set of distributions is restricted to the ones that can be fully factorized to simplify computations.
Then the log-likelihood of our model stated in (9) can be lower bounded as follows. Let $q_{\xi}(\bm{z})$ be the auxiliary variational distribution characterized by parameter $\bm{\xi}$ approximating the distribution $p_{\eta}(\bm{z})$, then the log-likelihood has a lower-bound, calculated as follows:
@@
$\geq$ $\sum_{\bm{z}\in\mathcal{Z}}q_{\xi}(\bm{z})\log\left[\frac{L_{0}\left(\bm{g},\bm{x},\bm{z};\bm{\alpha},\bm{\beta},\bm{\eta}\right)}{q_{\xi}(\bm{z})}\right]$
@@
However, this variational problem is usually intractable unless we impose more structure on the problem. In practice, researchers restrict the set $\mathcal{Q}$ to a smaller set of tractable distributions (Wainwright and Jordan, 2008; Mele and Zhu, forthcoming). In the case of the stochastic blockmodel, it is useful and intuitive to restrict $\mathcal{Q}$ to the set of multinomial distributions
@@
with $\bm{\xi}_{i}$ being the variational parameters. We collect the vectors of variational parameters in the matrix $\bm{\xi}$. This leads to a tractable lower bound that can be written in closed-form
$\ell_{B}(\bm{g},\bm{x},\bm{\alpha},\bm{\beta},\bm{\eta};\bm{\xi})$$\equiv$ $\sum_{\bm{z}\in\mathcal{Z}}q_{\bm{\xi}}(\bm{z})\log\left[\frac{L_{0}\left(\bm{g},\bm{x},\bm{z};\bm{\alpha},\bm{\beta},\bm{\eta}\right)}{q_{\bm{\xi}}(\bm{z})}\right]$
@@
$\log\,\pi_{ij,kl}(g_{ij},\bm{x})$$\coloneqq$ $g_{ij}\log\left[\frac{\exp\left[u_{ij,kl}(\bm{\alpha},\bm{\beta})+u_{ji,lk}(\bm{\alpha},\bm{\beta})\right]}{1+\exp\left[u_{ij,kl}(\bm{\alpha},\bm{\beta})+u_{ji,lk}(\bm{\alpha},\bm{\beta})\right]}\right]$
$+$ $(1-g_{ij})\log\left[\frac{1}{1+\exp\left[u_{ij,kl}(\bm{\alpha},\bm{\beta})+u_{ji,lk}(\bm{\alpha},\bm{\beta})\right]}\right]$
@@
$u_{ij,kl}(\bm{\alpha},\bm{\beta})=u(\bm{x}_{i},\bm{x}_{j},z_{ik}=z_{jl}=1,\bm{z};\bm{\alpha},\bm{\beta})$
Note that $g_{ii}\coloneqq 0$ for all $i$ by definition. The covariate information $\bm{x}$ encodes in our setting categorical information on the level of repositories. We, therefore, introduce for the pair of repositories $(i,j)$ the covariate vector $\bm{x}_{ij}=(x_{ij,1},\ldots,x_{ij,p})$ encoding in the $k$th entry whether the respective repositories match on the $k$th covariate. Observing (A.1.1), we can write $\pi_{kl}(g_{ij},\bm{x}_{ij},\bm{z})=\pi_{ij,kl}(g_{ij},\bm{x})$ by dropping the dependence on the particular pair of repositories and the parameters $\alpha$ and $\beta$.
Generally, we iteratively maximize (A.1.1) in terms of $\bm{\eta},\bm{\alpha},$ and $\bm{\beta}$ conditional on $\bm{\xi}$ and then the other way around. We denote the values of $\bm{\xi},\bm{\eta},\bm{\alpha},$ and $\bm{\beta}$ in the $t$th iteration by $\bm{\xi}^{(s)},\bm{\eta}^{(s)},\bm{\alpha}^{(s)},$ and $\bm{\beta}^{(s)}$ and the two steps comprising the Variational Expectation Maximization algorithm are:
@@
    $\ell_{B}(\bm{g},\bm{x},\bm{\alpha}^{(s+1)},\bm{\beta}^{(s+1)},\bm{\eta}^{(s+1)};\bm{\xi}^{(s)})\,\geq\,\ell_{B}(\bm{g},\bm{x},\bm{\alpha}^{(s)},\bm{\beta}^{(s)},\bm{\eta}^{(s)};\bm{\xi}^{(s)});$
@@
    $\ell_{B}(\bm{g},\bm{x},\bm{\alpha}^{(s+1)},\bm{\beta}^{(s+1)},\bm{\eta}^{(s+1)};\bm{\xi}^{(s+1)})\,\geq\,\ell_{B}(\bm{g},\bm{x},\bm{\alpha}^{(s+1)},\bm{\beta}^{(s+1)},\bm{\eta}^{(s+1)};\bm{\xi}^{(s)}).$
@@
Taking first order conditions with respect to each parameter implies the following closed-form update rules for $\bm{\eta}$, and $\pi_{kl}^{(s+1)}(d,x_{1},\ldots,x_{p},\bm{z})$ follow
$\eta_{k}^{(s+1)}:=\frac{1}{n}\sum_{i=1}^{n}\xi_{ik}^{(s+1)},\quad k=1,\ldots,K,$
@@
$\pi_{kl}^{(s+1)}(d,x_{1},\ldots,x_{p},\bm{z}):=\frac{\sum_{i=1}^{n}\sum_{j\neq i}\xi_{ik}^{(s+1)}\xi_{jl}^{(s+1)}\bm{1}\{g_{ij}=d,X_{1,ij}=x_{1},\ldots,X_{p,ij}=x_{p}\}}{\sum_{i=1}^{n}\sum_{j\neq i}\xi_{ik}^{(s+1)}\xi_{jl}^{(s+1)}\bm{1}\{X_{1,ij}=x_{1},\ldots,X_{p,ij}=x_{p}\}},$
for $k,l=1,\ldots,K$ and $d,x_{1},\ldots,x_{p}\in\{0,1\}$, respectively. The variables $X_{p,ij}\coloneqq\bm{1}\{x_{ip}=x_{jp}\}$ are indicators for homophily. Generalizations are possible, as long as we maintain the discrete nature of the covariates $\bm{x}$. Including continuous covariates is definitely possible, but it may result in a significant slowdown of this algorithm. We note that when running the Variational EM (VEM) algorithm, we are not interested in estimating the parameters $\bm{\alpha}$, $\bm{\beta}$ and $\gamma$, but only the estimation of probabilities $\pi_{kl}(g,\bm{x},\bm{z})$ for $g\in\{0,1\}$ and $\bm{x}\in\mathcal{X}$, where $\mathcal{X}$ is the set of all possible outcomes of the categorical covariates used in Step 1. This feature allows us to speed up the computation of several orders of magnitude (Dahbura et al., 2021).
@@
For very large networks, maximizing the lower bound in the second step with respect to $\bm{\xi}$ for given estimates of $\bm{\eta},\bm{\alpha},$ and $\bm{\beta}$ is still cumbersome and impractically slow. We thus borrow an idea from Vu et al. (2013) and find a minorizing function $M\left(\bm{\xi};\bm{g},\bm{x},\bm{\alpha},\bm{\beta},\bm{\eta},\bm{\xi}^{(s)}\right)$ for the lower bound. This consists of finding a function approximating the tractable lower bound $\ell_{B}(\bm{g},\bm{x},\bm{\alpha},\bm{\beta},\bm{\eta};\bm{\xi})$, but simpler to maximize. Such function $M\left(\bm{\xi};\bm{g},\bm{x},\bm{\alpha},\bm{\beta},\bm{\eta},\bm{\xi}^{(s)}\right)$ minorizes $\ell_{B}(\bm{g},\bm{x},\bm{\alpha},\bm{\beta},\bm{\eta};\bm{\xi})$ at parameter $\bm{\xi}^{(s)}$ and iteration $s$ of the variational EM algorithm if
$M\left(\bm{\xi};\bm{g},\bm{x},\bm{\alpha},\bm{\beta},\bm{\eta},\bm{\xi}^{(s)}\right)$$\leq$ $\ell_{B}(\bm{g},\bm{x},\bm{\alpha},\bm{\beta},\bm{\eta};\bm{\xi})\text{ \ \ for all }\bm{\xi}$

### v2

@@ 4.2 Results
Table 5: Parameter estimates and standard errors.
@@
Estimate Estimate
(Std. Error) (Std. Error)
Edges ($\alpha_{1}$) $-5.801$ $-7.979$
(0.276) (0.006)
$\log(n)\times$ Edges $\left(\alpha_{1}^{\ell}\right)$ $-0.533$
Externality ($\gamma$) $-0.095$
Maturity ($\beta_{1}$) 1.791 $-0.222$
(0.305) (0.011)
$\log(n)\times$ Maturity $\left(\beta_{1}^{\ell}\right)$ $-0.123$
Popularity ($\beta_{2}$) 1.944 0.118
(0.305) (0.01)
$\log(n)\times$Popularity $\left(\beta_{2}^{\ell}\right)$ $-0.18$
Size ($\beta_{3}$) 1.114 $-0.07$
(0.31) (0.011)
$\log(n)\times$Size$\left(\beta_{3}^{\ell}\right)$ $-0.122$
We estimate a model with $K=10$ unobserved types and initialize the types allocation for our algorithm using the Walktrap algorithm. [^24] As suggested by Wainwright and Jordan (2008) and others, we ran the algorithm with multiple initial guesses of the type allocations, but report the results that achieved the highest lower bound. While there is no established method to infer $K$ in this class of models, we have run the estimation with $K=5$ and $K=15$ for robustness, but in both cases the model’s fit was worse than with $K=10$.
The results of our estimation are reported in Table LABEL:tbl:results_python. The first column shows estimates of the structural parameters for links of the same (unobservable) type, while the remaining column provides the estimates for links of different types.
We estimate the externality ($\gamma$) to be negative, providing empirical evidence of the interdependence of links suggested by our strategic model. Essentially, our results provide evidence that when developers form connections, they consider the indirect impacts of their actions, at least partially. This behavior leads to a network architecture that might not be as efficient as one where a central planner aims to maximize the overall utility of all participants. Indeed, the externality encourages developers to establish fewer connections than would be ideal from an aggregate utility-maximization perspective. Maintainers see the number of dependencies of a package they consider linking to as a cost, thus leading to linking decisions that decrease the density of the aggregate network. The negative externality has further implications for contagion, as we explain in the next section.
The parameters $\alpha_{1}$ and $\alpha_{1}^{\ell}$ govern the density of the network and are both estimated negative. The parameter $\alpha_{1}$ governing links between types is likewise estimated negative. The magnitudes of the estimated parameters imply that—other things being equal—in equilibrium sub-networks of libraries of the same type tend to be less sparse than between-types. The sparsity increases with the number of libraries belonging to a particular type. One interpretation is that $\alpha$ measures costs of forming links, which is higher when there are more available libraries to choose from, possibly because of the need to monitor and evaluate more code when forming a decision.
By contrast, the parameters $\beta_{1},\beta_{2},\beta_{3}$ are all positive within types, indicating that developers are more likely to link to packages of the same type of similar maturity, popularity and size. On the other hand, this homophily seems attenuated by the size of the type, as shown by the negative coefficients of the parameters $\beta_{1}^{\ell},\beta_{2}^{\ell},\beta_{3}^{\ell}$.
When forming dependencies with packages of a different type, the effect of the observables is negative for maturity and popularity, while positive for size. We interpret this as evidence of complementarities across types.
One interpretation consistent with these results is that developers use popular features of same-type packages to include in their own code but use infrastructure-style code from different-type packages, which is not necessarily very popular.
4.3 Model Fit
To assess model fit, we follow the methodology developed in the literature on ERGM and HERGMs and simulate the model under the estimated parameters (Hunter et al., 2008; Babkin et al., 2020). We compute network statistics on those simulated networks and compare them with the corresponding statistics evaluated on the observed network. If the model fits the data, most of the simulated network statistics should resemble their observed counterparts.
In Figure 2 we show a boxplot with summaries of the networks obtained from 100 simulations of the model. The summaries show in Figure 2 are natural summaries, because they are the sufficient statistics of the model conditional on the types of packages.  [^25] We start the simulations at the observed network, and we use a Metropolis-Hastings sampler to update the network. We first run a burn-in of 500,000 iterations, and then sample a network every 50,000 iterations. These summaries are normalized, so that a value of zero indicates a perfect match. The red line is the observed network, so a perfect fit corresponds to all the network statistics concentrated around the red line. The boxplots show that the model is able to capture most of the aggregate structural properties of the Python dependency network: most of the simulated statistics are centered around the observed network. The only exception is the externality (i.e. the number of two-paths in our model), however the observed network falls in the $95\%$ confidence interval of the simulation output. We conclude that the estimated model does a good job in reproducing important features of the network.
Figure 2: Model fit: Network statistics
In Figure 3 we report similar boxplots for in-degree and out-degree, showing that the simulated networks replicate the observed one quite closely. These results are particularly striking as it is usually quite hard to fit such a model for a very large network.
Figure 3: Model Fit: Degree distributions
In the Online Appendix C, we compare the fit of our model to alternative approaches. An important question is whether the unobserved heterogeneity included in our model is needed. To answer this question we compare our estimated model to a model without unobserved types ($K=1$). This corresponds to a standard Exponential Random Graph model (ERGM), as studied in Mele (2017). The ERGM is not able to match the aggregate density of the network, while our model’s simulations are concentrated around the observed number of links. The second question is whether the externalities are important in explaining the topology of the equilibrium network. Therefore, we compare our model to a Stochastic blockmodel with covariates and $K=10$ types. This corresponds to our model with $\gamma=0$. Our simulations show that the our model does a better job matching the externality within types, and should therefore be preferred. Finally, while there is no currently established method to choose the optimal number of types for this class of complex models, we have done a significant effort in setting an appropriate $K$ for estimation. Our practical approach consists of setting the number of types by checking the model’s fit for different values of $K$. In appendix C we show the fit with $K=5$ and $K=15$. These alternative models overestimate or underestimate the externality level, so our model with $K=10$ is preferred.
@@ 5 Contagious Vulnerabilities
We adapt these ideas to the contagion of vulnerabilities in software dependency graphs. The prevalence of vulnerabilities in software motivated the creation of the Common Vulnerabilities and Exposure (CVE) program and the National Vulnerability Database (NVD) by the National Institute of Standards and Technology, which is an agency of the United States Department of Commerce (NIST, 2024). A software package depending on a vulnerable package is potentially vulnerable itself. To see how, consider, for example, the Equifax data breach (Federal Trade Commission, 2024) in 2017, caused by a vulnerability in the popular open source web application framework Apache Struts 2.262626Apache Struts 2 did not properly validate the Content-Type header of incoming http requests when processing file uploads, which allowed attackers to create http requests that include malicious code. Due to the lack of content validation, this allowed attackers to execute arbitrary code on the web server, ultimately granting them full control of the server. For details see the official NIST announcement on CVE-2017-5638 (https://nvd.nist.gov/vuln/detail/CVE-2017-5638). Equifax did not itself develop or maintain Apache Struts 2, but instead used it as part of its own code base. The vulnerability allowed attackers to execute malicious code providing access to the server hosting the web application developed by Equifax using Apache Struts 2 and costing the company more than \$1.5 Billion to date. [^27] See https://www.bankinfosecurity.com/equifaxs-data-breach-costs-hit-14-billion-a-12473. Article accessed 27 January 2024.
@@ 6 Conclusion
The network of dependencies among software packages is an interesting laboratory to study network formation and how incentives and externalities shape the topology of the network. Indeed, the creation of modern software gains efficiency by re-using existing libraries; on the other hand, dependencies expose new software packages to bugs and vulnerabilities from other libraries. This feature of the complex network of dependencies motivates the interest in understanding the incentives and equilibrium mechanisms driving the formation of such networks.
@@
Using data from the Python dependency network, we find evidence that coders create negative externalities for other coders when creating a link. This raises more questions about the formation of dependency graphs. Using our estimated model, we study the vulnerability of the dependency network to spread of a bug. We measure systemic risk as the average number of downstream packages that are affected by a vulnerability, assuming that the contagion follows a SIR model of epidemic. We simulate interventions that target nodes according to their in-degree, expected fatality and betweenness centrality. Such interventions are unlikely to significantly decrease the risk profile of the network. On the other hand, the increase in AI-assisted tools for programmers, may have a beneficial effect by reducing the cost of producing code in-house, without much need for dependencies.
While we have focused on a stationary realization of the network, there are important dynamic considerations in the creation of these dependencies. While the modeling of forward-looking maintainers may be useful to develop intuition about intertemporal strategic incentives and motives, we leave this development to future work, as it involves several complications in the estimation. Finally, we have focused on a single language, but the analysis can be extended to other languages as well, such as Java, R, Rust, C++ and others.
@@ Appendix A Computational details for estimation
where each package’s type is i.i.d. multinomial, so the distribution $p_{\bm{\eta}}(\bm{z})$ is
@@ A.1.1 Variational EM algorithm with MM updates
Then under some conditions – namely, that the network is large, each block/type is not too large with respect to the network, and the employed statistics are in some sense well-behaved (Babkin et al., 2020; Schweinberger, 2020; Schweinberger and Stewart, 2020) – we have
$L\left(\bm{g},\bm{x},\bm{z};\bm{\alpha},\bm{\beta},\gamma,\bm{\eta}\right)\approx L_{0}\left(\bm{g},\bm{x},\bm{z};\bm{\alpha},\bm{\beta},\bm{\eta}\right).$ (20)
Denoting by $q_{\xi}(\bm{z})$ a the auxiliary distribution characterized by parameter $\bm{\Xi}$ approximating the distribution $p_{\eta}(\bm{z})$, the log-likelihood of our model stated in (18) can be lower bounded as follows:
@@
$\geq$ $\sum_{\bm{z}\in\mathcal{Z}}q_{\xi}(\bm{z})\log\left(\frac{L_{0}\left(\bm{g},\bm{x},\bm{z};\bm{\alpha},\bm{\beta},\bm{\eta}\right)}{q_{\xi}(\bm{z})}\right)$
@@
However, this variational problem is usually intractable unless we impose more structure on the problem. In practice, researchers restrict the set $\mathcal{Q}$ to a smaller set of tractable distributions (Wainwright and Jordan, 2008; Mele and Zhu, 2023). In the case of the stochastic blockmodel, it is useful and intuitive to restrict $\mathcal{Q}$ to the set of multinomial distributions
@@
where we collect the variational parameters in the matrix $\bm{\Xi}=(\xi_{ik})\in[0,1]^{N\times K}$. This leads to the following tractable lower bound of $\ell$
$\ell_{B}(\bm{g},\bm{x},\bm{\alpha},\bm{\beta},\bm{\eta};\bm{\Xi})$$\equiv$ $\sum_{\bm{z}\in\mathcal{Z}}q_{\bm{\Xi}}(\bm{z})\log\left(\frac{L_{0}\left(\bm{g},\bm{x},\bm{z};\bm{\alpha},\bm{\beta},\bm{\eta}\right)}{q_{\bm{\Xi}}(\bm{z})}\right)$
@@
$\log\,\pi_{ij,kl}(g_{ij},\bm{x})$$\coloneqq$ $g_{ij}\log\left(\frac{\exp\left(u_{ij,kl}(\bm{\alpha},\bm{\beta})+u_{ji,lk}(\bm{\alpha},\bm{\beta})\right)}{1+\exp\left(u_{ij,kl}(\bm{\alpha},\bm{\beta})+u_{ji,lk}(\bm{\alpha},\bm{\beta})\right)}\right)$
$+$ $(1-g_{ij})\log\left(\frac{1}{1+\exp\left(u_{ij,kl}(\bm{\alpha},\bm{\beta})+u_{ji,lk}(\bm{\alpha},\bm{\beta})\right)}\right)$
@@
$u_{ij,kl}(\bm{\alpha},\bm{\beta})=u(\bm{x}_{i},\bm{x}_{j},z_{ik}=z_{jl}=1,\bm{z};\bm{\alpha},\bm{\beta}).$
Note that $g_{ii}\coloneqq 0$ for all $i$ by definition. The covariate information $\bm{x}$ encodes in our setting categorical information on the level of repositories (maturity, popularity, and size). We, therefore, introduce for the pair of repositories $(i,j)$ the covariate vector $\bm{x}_{ij}=(x_{ij,1},\ldots,x_{ij,p})$, where the $k$th entry is a binary indicator whether repositories $i$ and $j$ match on the $k$th covariate. The variables $x_{p,ij}\coloneqq\bm{1}\{x_{ip}=x_{jp}\}$ are, therefore, indicators for homophily. Generalizations are possible, as long as we maintain the discrete nature of the covariates $\bm{x}$. Including continuous covariates is possible, but this would result in a significant slowdown of this algorithm. Observing (A.1.1), we can drop the dependence on the particular pair of repositories and write $\pi_{kl}(g_{ij},\bm{x}_{ij})=\pi_{ij,kl}(g_{ij},\bm{x})$.
We maximize (A.1.1) in a blockwise manner,first in terms of $\bm{\eta},\bm{\alpha},$ and $\bm{\beta}$ conditional on $\bm{\Xi}$ and then the other way around. Denoting the values of $\bm{\Xi},\bm{\eta},\bm{\alpha},$ and $\bm{\beta}$ in the $t$th iteration by $\bm{\Xi}^{(s)},\bm{\eta}^{(s)},\bm{\alpha}^{(s)},$ and $\bm{\beta}^{(s)}$, the two steps comprising the Variational Expectation Maximization algorithm are:
@@
    $\ell_{B}\left(\bm{g},\bm{x},\bm{\alpha}^{(s+1)},\bm{\beta}^{(s+1)},\bm{\eta}^{(s+1)};\bm{\Xi}^{(s)}\right)\,\geq\,\ell_{B}\left(\bm{g},\bm{x},\bm{\alpha}^{(s)},\bm{\beta}^{(s)},\bm{\eta}^{(s)};\bm{\Xi}^{(s)}\right);$
@@
    $\ell_{B}\left(\bm{g},\bm{x},\bm{\alpha}^{(s+1)},\bm{\beta}^{(s+1)},\bm{\eta}^{(s+1)};\bm{\Xi}^{(s+1)}\right)\,\geq\,\ell_{B}\left(\bm{g},\bm{x},\bm{\alpha}^{(s+1)},\bm{\beta}^{(s+1)},\bm{\eta}^{(s+1)};\bm{\Xi}^{(s)}\right).$
@@
Taking first order conditions with respect to each parameter implies the following closed-form update rules for $\bm{\eta}$ and $\pi_{kl}^{(s+1)}(d,x_{1},\ldots,x_{p})$:
$\eta_{k}^{(s+1)}:=\frac{1}{n}\sum_{i=1}^{n}\xi_{ik}^{(s)},\quad k=1,\ldots,K,$ (25)
@@
$\pi_{kl}^{(s+1)}(d,x_{1},\ldots,x_{p}):=\frac{\sum_{i=1}^{n}\sum_{j\neq i}\xi_{ik}^{(s)}\xi_{jl}^{(s)}\bm{1}\{g_{ij}=d,X_{1,ij}=x_{1},\ldots,X_{p,ij}=x_{p}\}}{\sum_{i=1}^{n}\sum_{j\neq i}\xi_{ik}^{(s)}\xi_{jl}^{(s)}\bm{1}\{X_{1,ij}=x_{1},\ldots,X_{p,ij}=x_{p}\}},$ (26)
for $k,l=1,\ldots,K$ and $d,x_{1},\ldots,x_{p}\in\{0,1\}$, respectively. The probability $\pi_{kl}^{(s+1)}(d,x_{1},\ldots,x_{p})$ is the same for all packages $i$ and $j$ such that $z_{i}=k$ and $z_{j}=l$. Therefore, we Note that we are not interested in finding particular estimates for $\bm{\alpha}$, $\bm{\beta}$ and $\gamma$ when running the Variational EM (VEM) algorithm. We are only estimating the probabilities $\pi_{kl}(g,\bm{x})$ for $g\in\{0,1\}$ and $\bm{x}\in\mathcal{X}$, where $\mathcal{X}$ is the set of all possible outcomes of the categorical covariates used in Step 1. This feature allows us to speed up the computation of several orders of magnitude (Dahbura et al., 2021). Finally, remark that updates (25) and (26) can both be comprehended as functions of $\bm{\Xi}^{(s)}$.
@@
For very large networks, maximizing the lower bound in the second step with respect to $\bm{\Xi}$ for given estimates of $\bm{\eta},\bm{\alpha},$ and $\bm{\beta}$ is even under the variational approximation of (22) cumbersome and impractically slow. We thus borrow an idea from Vu et al. (2013) define a function $M\left(\bm{\Xi};\bm{g},\bm{x},\bm{\alpha},\bm{\beta},\bm{\eta},\bm{\Xi}^{(s)}\right)$ that minorizes the lower bound $\ell_{B}(\bm{g},\bm{x},\bm{\alpha},\bm{\beta},\bm{\eta};\bm{\Xi})$ in $\bm{\Xi}^{(s)}$. Therefore,
$M\left(\bm{\Xi};\bm{g},\bm{x},\bm{\alpha},\bm{\beta},\bm{\eta},\bm{\Xi}^{(s)}\right)$$\leq$ $\ell_{B}(\bm{g},\bm{x},\bm{\alpha},\bm{\beta},\bm{\eta};\bm{\Xi})\text{ \ \ for all }\bm{\Xi}\in[0,1]^{N\times K}$